# NOD-TAMP: Generalizable Long-Horizon Planning with Neural Object Descriptors

### v3

@@ Abstract
Keywords: Robot Learning, Robot Planning, Manipulation
@@ 3.2 Skill Representation
We employ NDFs in our skills to represent not only their control trajectories but also their start and end states. Accordingly, we represent each skill as a tuple: . Here, name denotes the skill type (e.g., Pick, Insert). are skill parameters, which include the skill-relevant objects and their observed point clouds. Preconditions pre and effects eff specify the constraints that must hold before skill execution and the resulting new constraints, respectively. A constraint is represented by the relative configuration of two point clouds. For example, an Insert skill may require the robot to hold a object in a specific way. Executing the skill results in a new constraint between the object and a receptacle. Finally, let be a set of end-effector poses for this skill. A core objective of our method is to compose coherent multi-step plans by selecting and adapting a suitable trajectory within each constituent skill. We include details of all skills used in our experiments in the supplementary material.
@@ 4.1 Skill Adaptation
We seek to adapt a skill to a newly observed scene, which may be populated with new objects and layouts. To do so, we leverage a key invariance: the skill trajectory still needs to satisfy the recorded constraints (i.e., relative configurations between pairs of objects). Our skill adaptation module (1) transforms the skill trajectories to constraint-centric NDF feature trajectories and (2) adapts the trajectory to the observed scene via sequential optimization. To encode a skill trajectory relative to a recorded constraint, we consider common rearrangement skills that can be divided into two categories: hand-object interaction, such as grasping and manipulating constrained mechanisms (doors, etc.), and object-object interaction, where a robot uses the object in hand to interact with another object, such as placing and insertion.
@@ 4.2 Skill Planning
For a given task, we assume an -step task plan skeleton that defines a sequence of selected skills, e.g., [Pick(mug), Place(mug, bin), …]. Recall that each skill can contain multiple candidate demonstration trajectories. The start and end of each trajectory represent its precondition and effect constraint, respectively. The essential step in skill planning is, for each skill in the plan, choose a candidate trajectory that is most compatible with the constraints of its adjacent skills. We calculate compatibility based on the distance between pairs of constraints in the NDF space. For simplicity, for the -th skill in the plan, we denote the NDF-encoded precondition of a candidate skill trajectory as , and the effect as . The compatibility is calculated as Finally, we parse the goal configuration into a set of pair-wise object constraints and encode them as a set of NDF features . We then compute the plan cost as the distance between and the final accumulated constraints of the entire plan sequence.
After we obtain all costs for all skill trajectory combinations, the plan with the lowest plan-wide NDF feature distance is returned. For simplicity, we present this as a Cartesian product over relevant skills, but this can be done more efficiently by performing a Uniform Cost Search in plan space, where the NDF feature distance serves as a the cost function. Algorithm 2 in the supplementary material displays the pseudocode for the NOD-TAMP planner.
@@ 5 Experiments
We validate NOD-TAMP and how its components contribute to solve long-horizon tasks, perform fine-grained manipulation, and generalize to new object shapes. We select three evaluation settings: (1) LIBERO [15], a standard manipulation benchmark that feature diverse objects and long-horizon tasks, (2) a set of custom tabletop tasks that stress test spatial generalization and skill reasoning&reuse, and (3) six real-world tasks with noisy perception and multitudes of challenges combined. We highlight key conclusions in this section and leave additional results and experiment detail in the supplementary material.
@@ 5.3 Evaluation on the LIBERO Benchmark
We hypothesize that the performance gap is caused by different state/action representations and the structural information leveraged to build the system. The BC methods directly learn a mapping from the scene observation to the actions and thus require a huge broad data to cover diverse situations. Our method, along with the action-transferring baselines (e.g., NDF, MimicGen), utilize object-centric representations to adapt the spatial correspondences from demo scenes to test scenes. MimicGen+ assumes identical correspondences between demonstration objects and test objects, directly replaying trajectories in the local frames of test objects. In contrast, our approach leverages learned object representations to infer spatial correspondences, enabling the transferred actions to be more robust to variations in object geometry. Compared to NDF+, NOD-TAMP transfers a sequence of actions that represent each dynamic manipulation skill instead of just a single target state, in this case, the last end-effector pose in the demonstration trajectory. This improves NOD-TAMP’s performance in fine-grained manipulation tasks.
@@ 5.4 Evaluation on Customized Tabletop Tasks
NOD-TAMP is able to solve low-tolerance manipulation tasks. The ToolHang task requires fine-grained manipulation skills that adapt to various object poses. NOD-TAMP achieves 75% success rate, outperforming NDF+ and MimicGen+ that cannot adapt their trajectories based on environment changes. Since the task does not require avoiding obstacles and reasoning over grasp poses, we find that ablated baselines (Ours/MP and Ours/SR) achieve similar performance as our full method.
@@ 5.5 Real-world Evaluation
6 Limitations
A limitation of NOD-TAMP is the computation efficiency in NOD-based trajectory adaptation (more detailed analysis in supplementary). We hypothesize that this bottleneck can be addressed via lightweight neural networks or more efficient optimization techniques, which is beyond the scope of this work. Additionally, we are interested in incorporating TAMP constraints that are parameterized by NOD features to, for example, ensure that a mug containing liquid is always upright in a plan.
7 Conclusions
@@ Appendix A Table of Contents
Real-World Experiments (Sec. B): Elaborate on our real-world experimental setting and analyze the results;
@@ Appendix E Robustness to Perception Noise
To understand the performance of our system under different levels of perception noise, such as levels present in real-world sensors, we perform an experiment where we inject noise in the point cloud observation in simulation. We peform evaluation on the first stage of the ToolHang task, a high-precision task with tolerance of approximately 5mm. The robot needs to pick up the frame object and insert it into a stand. A study of the depth accuracy of the Microsoft Azure Kinect [57] showed that, within a distance of meters, the noise standard deviation is meters.

### v4

@@ Abstract
$\ast$$\ast$footnotetext: Equal Contribution.Keywords: Robot Learning, Robot Planning, Manipulation
@@ 3.2 Skill Representation
We employ NDFs in our skills to represent not only their control trajectories but also their start and end states. Accordingly, we represent each skill as a tuple: . We show a concrete example of skill representation in supplementary material Fig. 16. Here, name denotes the skill type (e.g., Pick, Insert). param } are skill parameters, which include the skill-relevant objects and their observed point clouds. Preconditions pre specifies a set of constraints that must hold before skill execution. The eff is change in conditions, i.e., adding or removing constraints, as a result of successfully executing the skill. A constraint is represented by the relative configuration of two point clouds, which is encoded through NDF features (e.g., ). For example, an Insert skill may require the robot to hold a object in a specific way. Executing the skill results in a new constraint between the object and a receptacle. Finally, let be a set of end-effector poses for this skill. A core objective of our method is to compose coherent multi-step plans by selecting and adapting a suitable trajectory within each constituent skill. We include details of all skills in our experiments in the supplementary material.
@@ 4.1 Skill Adaptation
We seek to adapt a skill to a newly observed scene, which may be populated with new objects and layouts. To do so, we leverage a key invariance: the skill trajectory still needs to satisfy the recorded constraints (i.e., relative configurations between pairs of objects). Our skill adaptation module (1) transforms the skill trajectories to constraint-centric NDF feature trajectories using Equation 3.1 and (2) adapts the trajectory to the observed scene via sequential optimization using Equation 3.1. Specifically, to encode a recorded skill trajectory and adapt it to the test scenarios, we consider common rearrangement skills that can be divided into two categories: hand-object interaction, such as grasping and manipulating constrained mechanisms (doors, etc.), and object-object interaction, where a robot uses the object in hand to interact with another object, such as placing and insertion.
@@ 4.2 Skill Planning
While our system can generate task plans independently or be embedded in an outer TAMP algorithm thanks to our skills functioning as PDDL operators, we assume the task plan is given. Our focus is on the bi-level planning problem of chaining and adapting skill trajectories to create coherent long-horizon motion plans. For a given task, we assume an -step task plan skeleton that defines a sequence of selected skills, e.g., [Pick(mug), Place(mug, bin), …]. Recall that each skill can contain multiple candidate demonstration trajectories. The start and end of each trajectory represent its precondition and effect constraint, respectively. The essential step in skill planning is, for each skill in the plan, choose a candidate trajectory that is most compatible with the constraints of its adjacent skills. We calculate compatibility based on the distance between pairs of constraints in the NDF space. For simplicity, for the -th skill in the plan, we denote the NDF-encoded precondition of a candidate skill trajectory as , and the effect as . The compatibility is calculated as Finally, we parse the goal configuration into a set of pair-wise object constraints and encode them as a set of NDF features . We then compute the plan cost as the distance between and the final accumulated constraints of the entire plan sequence.
We use Depth-first search (DFS) to traverse over possible skill plans. After we obtain all costs for all skill trajectory combinations, the plan with the lowest plan-wide NDF feature distance is returned. For simplicity, we present this as a Cartesian product over relevant skills, but this can be done more efficiently by performing a Uniform Cost Search in plan space, where the NDF feature distance serves as a the cost function. Algorithm 2 in the supplementary material displays the pseudocode for the NOD-TAMP planner.
@@ 5 Experiments
We validate NOD-TAMP and how its components contribute to solve long-horizon tasks, perform fine-grained manipulation, and generalize to new object shapes. We select three evaluation settings: (1) LIBERO [15], a standard manipulation benchmark that feature diverse objects and long-horizon tasks, (2) a set of custom tabletop tasks that stress test spatial generalization and skill reasoning&reuse, and (3) six real-world tasks with noisy perception and multitudes of challenges combined. We highlight key conclusions here and leave additional details in the supplementary material.
@@ 5.3 Evaluation on the LIBERO Benchmark
We hypothesize that the performance gap is caused by different state/action representations and the structural information leveraged to build the system. The BC methods directly learn a mapping from the scene observation to the actions and thus require a huge broad data to cover diverse situations. Our method, along with the action-transferring baselines (e.g., NDF, MimicGen), utilize object-centric representations to adapt the spatial correspondences from demo scenes to test scenes. MimicGen+ assumes identical correspondences between demonstration objects and test objects, directly replaying trajectories in the local frames of test objects. In contrast, our approach leverages learned object representations to infer spatial correspondences, enabling the transferred actions to be more robust to variations in object geometry. Compared to NDF+, NOD-TAMP transfers a sequence of actions that represent each dynamic manipulation skill instead of just a single target state, in this case, the last end-effector pose in the demonstration trajectory. This improves NOD-TAMP’s performance in fine-grained manipulation tasks. As LIBERO tasks are less constrained on grasping strategies (e.g., grasping the bowl to place it on a plate), we found the O/SR baseline performs similarly to our full system.
@@ 5.4 Evaluation on Customized Tabletop Tasks
NOD-TAMP is able to solve low-tolerance manipulation tasks. The ToolHang task requires fine-grained manipulation skills that adapt to various object poses. NOD-TAMP achieves 75% success rate, outperforming NDF+ and MimicGen+ that cannot adapt their trajectories based on environment changes. Since this task does not require obstacle avoidance or constrained grasping strategies, our ablated baselines (Ours/MP, Ours/SR, and NSC) perform similarly to our full method.
@@ 5.5 Real-world Evaluation
6 Conclusions and Limitations
@@
A limitation of NOD-TAMP is the computational efficiency in NOD-based trajectory adaptation. This bottleneck might be addressed with lightweight neural networks or more efficient optimization techniques. For skill planning, we use DFS to traverse over possible skill plans, this part can be further accelerated through optimization in the implementation (e.g., parallel DFS or building K-D trees in the NDF constraint space) to accommodate large-scale demos. Besides, extending the system to handle deformable objects would necessitate adapting the current NOD to account for time-varying features, which represents a challenging but interesting area for future work.
7 Acknowledgment
This work was partially supported by NSF Award #2409016 and NSF Award #2101250.
@@ Appendix A Table of Contents
Real-World Experiments (Sec. B): Elaborate on our real-world experimental setting and analyze the results
@@ Appendix E Robustness to Perception Noise
To understand the performance of our system under different levels of perception noise, such as levels present in real-world sensors, we perform an experiment where we inject noise in the point cloud observation in simulation. We peform evaluation on the first stage of the “ToolHang” task, a high-precision task with tolerance of approximately 5mm. The robot needs to pick up the frame object and insert it into a stand. A study of the depth accuracy of the Microsoft Azure Kinect [57] showed that, within a distance of meters, the noise standard deviation is meters.